### v1

# Application of Quantum Annealing to Computation of Molecular Properties

Abstract
We present the results of our quantum annealing computations of the permanent electric dipole moments of several molecules. By applying an electric field as a perturbation and measuring the corresponding energy responses, the molecular electric dipole moments are obtained numerically through the finite field method. The ground-state electronic wavefunctions and energies are obtained using the quantum annealer eignsolver algorithm. This work provides a pathway for the computation of molecular properties in the quantum annealing paradigm.
Quantum annealing (QA) is a heuristic approach for solving combinatorial optimization problems that leverages quantum effects such as tunneling and superposition [1]. It represents a specialized approach closely related to Adiabatic Quantum Computing, designed to address computationally hard problems by harnessing these quantum effects [2]. The Quantum Annealer Eigensolver (QAE), a quantum-classical hybrid approach based on the variational principle [3], has been demonstrated to yield highly accurate results for ground-state energy optimization [4]. Accurate calculation of physical properties is crucial in chemistry and physics as it provides insight to fundamental properties of molecules and materials. This method is broadly applicable to a wide range of physical optimization problems, including molecular vibrational spectra [5], complex eigenvalue problems, energy calculations for molecular electronic states, relativistic fine-structure splitting in highly charged atomic ions [4], particle physics simulations [6], and lattice gauge theories [7].
In this work, we extend the application of QA to compute an important molecular property- the electric permanent dipole moment (PDM). This advancement not only underscores the broad applicability of the method but also opens new avenues for analyzing molecular properties beyond ground-state energy evaluation. We compute the PDM of molecules adapting the finite-field method (FFM) to QA. Moreover, PDMs are important for various applications, particularly for SrF [8], the first molecule to be laser-cooled [9] and and BaF, a promising candidate for precision measurements in fundamental physics [10].
This paper begins with an overview of the FFM [11] and the Dirac-Coulomb (DC) framework [12], which form the theoretical basis for relativistic calculations using quantum annealing. We then describe the methodology used in the Dirac–Fock (DF) calculations for the BeF, MgF, CaF, SrF, and BaF molecules. Following this, we introduce the QAE algorithm and explain its implementation on the D-Wave quantum annealer. Subsequently, the overall workflow of the problem is outlined. A comparative analysis of the results obtained from simulated annealing, quantum annealing, and full configuration interaction (FCI)/exact diagonalization(ED) is subsequently presented. The paper concludes with a summary of our key findings and a discussion of potential directions for future research.
[…]
$\langle\hat{O}\rangle\approx\lim_{\epsilon\to 0}\frac{E(+\epsilon)-E(-\epsilon)}{2\epsilon}.$ (1)
A small perturbative electric field –typically on the order of $10^{-4}$ to $10^{-3}$ a.u. is applied along the $z$-axis, and the total energy is computed for both positive and negative field values. Both $E(+\epsilon)$ and $E(-\epsilon)$ are computed using the QAE. The Hamiltonian that employed in this work is the DC Hamiltonian, $H_{DC}$, which is:
$H_{DC}=\sum_{i}\left[c{\alpha}\cdot{p}_{i}+\beta mc^{2}-\sum_{A}\frac{Z_{A}}{|{r}_{i}-{R}_{A}|}\right]+\sum_{i\neq j}\frac{1}{|{r}_{i}-{r}_{j}|}.$ (2)
Here, $c$ is the speed of light, $\alpha$} and $\beta$ are the Dirac matrices, and ${p}_{i}$ refers to the momentum of the $i^{th}$ electron. The summation over the electronic coordinates is denoted by $i$, and that over the nuclear coordinates is indicated by $A$. ${r}_{i}$ is the position vector from the origin to the site of an electron, and ${R}_{A}$ is the position vector from the origin to the coordinate of a nucleus. $Z_{A}$ is the atomic number of the $A^{th}$ nucleus. Then the second-quantized form of the Hamiltonian is given by
$H=\sum_{pq}h_{pq}a_{p}^{\dagger}a_{q}+\frac{1}{2}\sum_{pqrs}g_{pqrs}a_{p}^{\dagger}a_{q}^{\dagger}a_{s}a_{r}.$ (3)
[…]
The molecular PDMs in this study were first computed using classical computers to provide benchmarks for the values obtained via the QAE algorithm. Specifically, DF and relativistic coupled-cluster singles and doubles (RCCSD) calculations were performed using the DIRAC22 software package [14]. The DF computations were carried out at the self–consistent field (SCF) level, with a perturbative electric field ($\epsilon=\pm 10^{-3}$) applied along the $z$-axis to evaluate the field-dependent energies. To enhance computational efficiency, molecular symmetry was exploited through use of C2v double group symmetry in the molecular setup.
In addition to DF, RCCSD calculations were performed using the same code, employing the one- and two-electron integrals generated internally by DIRAC22. To enable a direct comparison with the QAE algorithm, the FCI Hamiltonian was constructed classically by applying the Slater–Condon rules [15] and subsequently diagonalized exactly to obtain the FCI/ED energy. This Hamiltonian was then mapped to the D-Wave quantum annealer, where the lowest eigenvalue corresponding to the ground-state energy was extracted. The ground-state energies corresponding to both positive and negative perturbation strengths ($\epsilon=\pm 10^{-3}$) were computed to evaluate the field-dependent energy shifts. QA was employed to determine the electronic energies within selected active spaces—specifically, $(8o,3e)$ and $(14o,7e)$ configurations. These electronic energies were then combined with the corresponding core energy and nuclear repulsion energy terms to obtain the total molecular energy. Using this total energy, the electronic component of the dipole moment was evaluated through the finite-field approach explicitly linked Eq.(1).
To obtain the PDM, the nuclear contribution was added based on experimental available bond lengths. The bond lengths used in the calculations were 1.361 Å for BeF, 1.75 Å for MgF, 1.967 Å for CaF, 2.075 Å for SrF, and 2.16 Å for BaF [16, 17, 18]. For all calculations, uncontracted Gaussian-type orbitals (GTOs) [19] were used, and the kinetic balance condition was strictly maintained to ensure the accuracy of relativistic treatments. The basis sets for the heavier atoms Sr and Ba were chosen from the dyall.c2v, dyall.c3v [20], and dyall.c4v [21, 19]. For lighter elements such as Be, Mg, Ca, and F, the exponential parameters from the correlation-consistent polarized valence (cc-pV) basis sets were used, obtained from the EMSL Basis Set Exchange Library [20].
Table 1: Atomic basis sets used in the calculations. This table lists the uncontracted GTOs employed for each atom. For Be, Mg, Ca, and F, correlation-consistent polarized valence (cc-pV) basis sets of increasing size were used. For the heavier Sr and Ba atoms, Dyall’s relativistic basis sets (dyall.c2v, dyall.c3v, dyall.c4v) were utilized.
   Atom Basis
Be cc-pVDZ: 9s, 4p, 1d
cc-pVTZ: 11s, 5p, 2d, 1f
cc-pVQZ: 12s, 6p, 3d, 2f, 1g
Mg cc-pVDZ: 12s, 8p, 1d
cc-pVTZ: 15s, 10p, 2d, 1f
cc-pVQZ: 16s, 12p, 3d, 2f, 1g
Ca cc-pVDZ: 14s, 11p, 5d
cc-pVTZ: 20s, 14p, 6d, 1f
cc-pVQZ: 22s, 16p, 7d, 2f, 1g
Sr dyall.c2v: 20s, 14p, 9d
dyall.c3v: 28s, 20p, 13d, 2f
dyall.c4v: 33s, 25p, 15d, 4f, 2g
Ba dyall.c2v: 25s, 19p, 13d
dyall.c3v: 31s, 25p, 15d, 2f
dyall.c4v: 36s, 30p, 18d
F cc-pVDZ: 9s, 4p, 1d
cc-pVTZ: 10s, 5p, 2d, 1f
cc-pVQZ: 12s, 6p, 3d, 2f, 1g
QA involves preparing the ground-state of an initial Hamiltonian, $H_{I}$, which is typically chosen to be the transverse field Hamiltonian, and then adiabatically transforming it into another Hamiltonian, $H_{F}$, generally chosen to be the Ising Hamiltonian whose ground-state encodes the solution to the problem. The transformation follows:
$H(t)=\left(1-\frac{t}{T}\right)H_{I}+\frac{t}{T}H_{F};\quad t\in[0,T],$ (4)
[…]
$\xi(\vec{c},\lambda)=\sum_{i,j=0}^{N-1}c_{i}c_{j}(H_{ij}-\lambda\delta_{ij}).$ (5)
[…]
$c_{i}=\sum_{\alpha=1}^{K-1}2^{\alpha-K}q_{\alpha}^{i}-q_{K}^{i}.$ (6)
This converts the problem from a continuous optimization over $\{c_{i}\}$ to a discrete optimization over the binary variables $\{q^{i}_{\alpha}\}$. Here, the first K1K-1 qubits encode the fractional part, the last qubit $q_{K}^{i}$ encodes the sign and So the each $c_{i}$ is repesented by $K$ number of binary variable $q^{i}_{\alpha}$ [24, 25, 5].
$\begin{split}\xi(\vec{q},\lambda)=\sum_{i,j=1}^{N}\sum_{\alpha,\beta=1}^{K-1}&\left(2^{\alpha+\beta-2K}q_{\alpha}^{i}q_{\beta}^{j}-2^{\alpha+1-K}q_{\alpha}^{i}+q_{K}^{i}q_{K}^{j}\right)\\ &\times\left(H_{ij}-\lambda\delta_{ij}\right).\end{split}$ (7)
The proposed mapping approach necessitates an expansion of the Quadratic Unconstrained Binary Optimization (QUBO) formulation [26] from dimensions $(N\times N)$ to $(NK\times NK)$. This dimensional increase is required to overcome the restricted connectivity limitations inherent in D-Wave’s quantum annealing architecture. So, the optimization of $\xi(\vec{q},\lambda)$ is implemented through a pipelined framework consisting of sequential sub-QUBO optimizations, where each sub-problem operates on an independent subset of coefficients. In our methodology, we first sort the coefficients in descending order basis on the basis of their classical values obtained from perturbation theory. These ordered coefficients are then partitioned into systematic groupings that form the basis for sub-QUBOs construction. These sub-QUBOs are then independently submitted to the annealer to obtain corresponding solution [4].
The embedding process maps the each sub-QUBO onto the D-Wave hardware’s native topology [27, 28] by forming chains of physical qubits to represent single logical variables. We used the default chain coupling strength to maintain consistent correlations during annealing while preventing chain dynamics from dominating the system’s energy landscape [29]. During the annealing process, we first map the subproblem onto the Pegasus graph structure of the D-Wave Advantage quantum processing unit. For each execution, we acquire a fixed ensemble of 1000 samples(anneals or number of reads), where each sample represents a specific configuration of the coefficients in the qubit state space.
After obtaining the binary solutions from each independent sub-QUBO problem from annealer, we aggregate those solutions and proceed to a post-processing stage. This stage involves applying a steepest descent optimization at every value of $\lambda$ within the defined search interval $\lambda_{-},\lambda_{+}]$. At every value of $\lambda$ the annealer returns binary values of $\vec{q}$}, totaling $NK$ bits. In our case, we have set $K=10$ to encode each parameter with sufficient resolution. The entire implementation utilizes the tools available in D-Wave’s Ocean SDK [30], and specifically employs the D-Wave Advantage QPU for execution.
This method begins with an $NK$-variable QUBO problem, due to the computational complexity of directly solving such large-scale problems, we employ a perturbative approximation to reduce the problem size while preserving the essential features of the solution space. In this work, the strategic partitioning of a large-scale QUBO problem into $M$ smaller, more tractable sub-QUBOs is employed. Each resulting sub-problem is then optimized independently to find a local solution, denoted as $\vec{q}^{(m)}$ for $m\in\{1,\dots,M\}$. The optimization of these sub-QUBOs can be performed using a variety of solvers, including quantum annealers or classical algorithms such as Simulated Annealing(SA). Although the decomposition methodology permits parallel execution on multiple quantum annealers, in this work the sub-QUBOs are optimized sequentially due to limited processor access, and thus the overall computational demand remains unchanged.
Figure 1: Workflow of the standard QAE algorithm, where an energy functional with Lagrange multiplier $\lambda$ is mapped to a QUBO, optimized on the quantum annealer, and refined via classical post-processing across a range of $\lambda$ values to obtain the minimum ground-state energy.
The global solution is constructed by combining partial solutions into a candidate vector $\vec{q}$. We employ steepest descent minimization within the energy landscape $\xi(\vec{q},\lambda)$, where at each iteration the algorithm selects the direction of maximal energy reduction. This refinement stage compensates for decomposition approximations by explicitly incorporating all intervariable couplings from the original Hamiltonian. The converged solution $\vec{q}_{\text{final}}$ attains accuracy comparable to direct methods.
Algorithm 1 QA for Hamiltonian Minimization
1:Full Hamiltonian matrix $H$ of size $N\times N$.
2:Initial guess for the parameter $\lambda\in[\lambda_{\text{min}},\lambda_{\text{max}}]$.
3:Optimal solution vector $\vec{c}_{\text{final}}$ with minimal energy.
4:Phase 1: Sub-QUBO Formulation
5:Partition the Hamiltonian $H$ into $M$ sub-QUBO problems, denoted as $\xi_{m}(\vec{q}^{(m)},\lambda)$ for $m\in\{1,\dots,M\}$.
6:Phase 2: Quantum Annealing Iteration
7:for $\lambda$ from $\lambda_{\text{min}}$ to $\lambda_{\text{max}}$ do $\triangleright$ Iterate over the range of $\lambda$
8:  for each sub-QUBO $\xi_{m}(\vec{q}^{(m)},\lambda)$ do
9:   Encode the classical problem into the sub-QUBO: $\xi_{m}(\vec{q}^{(m)},\lambda)\leftarrow\xi_{m}(\vec{c}^{(m)},\lambda)$.
10:   Perform quantum annealing on $\xi_{m}$: $q_{\text{raw}}^{(m)}\leftarrow\min[\xi_{m}(\vec{q},\lambda)]$ $\triangleright$ Find the optimal binary configuration for the $i$-th sub-QUBO.
11:  end for
12:  Combine the optimal solutions from all sub-QUBOs: $\vec{q}_{\text{raw}}\leftarrow\bigcup_{m=1}^{M}q_{\text{opt}}^{(m)}$.
13:Phase 3: Post-processing and Energy Evaluation
14:  Apply a steepest descent method for refinement: $\vec{q}_{\text{post-processed}}\leftarrow\text{SteepestDescent}(\vec{q}_{\text{raw}})$.
15:  Reverse the encoding to obtain classical coefficients: $\vec{c}_{\text{post-processed}}\leftarrow\text{ReverseEncoding}(\vec{q}_{\text{post-processed}})$.
16:  Compute the total energy $E_{\lambda}$ using the obtained coefficients: $E_{\lambda}\leftarrow\text{ComputeEnergy}(\vec{c}_{\text{post-processed}})$.
17:end for
18:Final Solution Selection
19:Identify the best solution: $\vec{c}_{\text{final}}\leftarrow\underset{\lambda\in[\lambda_{\text{min}},\lambda_{\text{max}}]}{\operatorname{argmin}}\{E_{\lambda}\}$. $\triangleright$ Select the coefficient vector corresponding to the minimum energy found.
20:Return $\vec{c}_{\text{final}}$
In Table 2 and 3, we present our computed PDM values using the QAE and benchmark them against conventional methods such as RCCSD, FCI, and SA results [31], all within two different active spaces but employing different basis sets as described in Table 1. The QAE computations were performed on both classical (SA) platforms [31] and quantum annealing hardware (D-Wave Advantage). These simulations explicitly take into account qubit connectivity constraints. The results indicate that QAE can accurately predict both total energies (for both perturbation $\pm\epsilon$)and PDM values, with deviations within acceptable limits. Specifically, the hardware-executed QAE results were found to be consistent with their SA and FCI counterparts as given in Table 4.
For the QAE calculations, we selected energy ranges $\lambda_{-},\lambda_{+}]$ slightly below and above the DF energy, corresponding to the active spaces $(8o,3e)$ and $(14o,7e)$ used in this study. After generating the full configuration interaction determinant basis, we employed a perturbative selection procedure to retain only the most relevant configurations. This step mitigates the exponential growth in the number of Slater determinants with increasing numbers of virtual orbitals. The Hamiltonian matrix was then constructed within the reduced determinant space and used for subsequent QAE calculations. For each specified energy range $\lambda_{-},\lambda_{+}]$, we performed three independent QAE runs which called repeats and selected the lowest energy [32, 5]. Additionally, sub-QUBOs were restricted to a maximum of 60 binary variables for practical implementation on the hardware in this calculation.
We now discuss possible sources of error in our calculation [29, 33]. These include hardware-induced errors, chain-breaking artifacts, and approximations made during sub-QUBOs generation [34]. In particular, during sub-QUBOs construction, certain quadratic interaction terms may be neglected, potentially impacting the final energy estimates. We also performed the same computations using simulated annealing on the classical processor with the same sub-QUBO settings and observed a negligible error, suggesting strong agreement between quantum and classical annealing techniques. Assuming that SA and ideal QA yield similar outcomes, we conclude that hardware noise likely has a minimal impact on the accuracy of PDMs. This stability may result from the use of a large number of anneals (e.g., 1000), which effectively average out random fluctuations and reduce noise effects.
Furthermore, we conducted simulations assuming all-to-all connectivity and using the same sub-QUBO decomposition as in the quantum hardware setup. This residual discrepancy is primarily attributed to limitations in our QUBO encoding scheme—specifically, the choice of $K=10$ in the fixed-point representation of the CI coefficients, yielding a precision of $\frac{1}{2^{10}}$ in energy. Increasing $K$ would improve numerical precision but also significantly enlarge the QUBO, rendering it less feasible for hardware execution at present. The most prominent deviation was observed between simulations assuming all-to-all connectivity [35] and those constrained by the native connectivity of the D-Wave hardware, indicating that chain formation and embedding limitations are key sources of inaccuracy [29].
Finally, we note that our method does not suffer from barren plateaus, which are common in approaches like VQE [36] due to the structure of the ansatz. In contrast, QAE directly optimizes the CI coefficients, and the energy remains sensitive to changes in these values. Since our approach is based on complete active space CI, it avoids the flat regions in the energy landscape that make optimization difficult. Moreover, we were able to successfully calculate properties such as the PDM using this method, demonstrating its robustness compared to VQE.
Table 2: Comparison of the PDM values, in debye, for the $(8o,3e)$ and $(14o,7e)$ active space. Results from the QAE are benchmarked against classical RCCSD, FCI, and SA methods for the molecules and basis sets studied.
Molecule Basis   RCCSD   FCI   SA   QAE
BeF $c$-pVDZ 1.338 1.342 1.351 1.354
$c$-pVTZ 1.242 1.245 1.239 1.267
$c$-pVQZ 1.270 1.271 1.273 1.267
MgF $c$-pVDZ 3.188 3.185 3.189 3.178
$c$-pVTZ 3.073 3.045 3.114 3.074
$c$-pVQZ 3.099 3.099 3.098 3.108
CaF $c$-pVDZ 2.685 2.688 2.686 2.684
$c$-pVTZ 2.687 2.685 2.686 2.688
$c$-pVQZ 2.684 2.683 2.680 2.681
SrF dyall.v2z 2.753 2.753 2.752 2.752
dyall.v3z 2.723 2.723 2.721 2.719
dyall.v4z 2.759 2.759 2.763 2.764
BaF dyall.v2z 2.323 2.323 2.323 2.322
dyall.v3z 2.222 2.222 2.222 2.222
dyall.v4z 2.297 2.297 2.296 2.293
Table 3: Comparison of the PDM values, in debye, for the $(14o,7e)$ active space. Results from the QAE are benchmarked against classical RCCSD, FCI, and SA methods for the molecules and basis sets studied.
Molecule Basis   RCCSD   FCI   SA   QAE
BeF $c$-pVDZ 1.318 0.421 0.471 0.635
$c$-pVTZ 1.318 0.418 0.448 0.914
$c$-pVQZ 1.268 1.268 1.275 1.267
MgF $c$-pVDZ 3.175 3.222 3.160 3.218
$c$-pVTZ 3.072 3.072 3.084 3.078
$c$-pVQZ 3.099 3.105 3.109 3.103
CaF $c$-pVDZ 2.700 2.686 2.712 2.654
$c$-pVTZ 2.694 2.688 2.685 2.695
$c$-pVQZ 2.690 2.684 2.681 2.680
SrF dyall.v2z 2.764 2.759 2.750 2.755
dyall.v3z 2.734 2.730 2.727 2.770
dyall.v4z 2.764 2.762 2.762 2.724
BaF dyall.v2z 2.324 2.322 2.315 2.322
dyall.v3z 2.225 2.223 2.224 2.219
dyall.v4z 2.299 2.298 2.296 2.298
Table 4: Percentage error for the PDM values calculated via the QAE and SA. The error is calculated relative to the benchmark FCI results presented in Tables II and III for the $(8o,3e)$ and $(14o,7e)$ active spaces, respectively
Molecule Basis (8o,3e) (14o,7e)
$Err_{QAE}$ $Err_{SA}$ $Err_{QAE}$ $Err_{SA}$
BeF $c$-pVDZ 0.894 0.693 50.837 11.738
$c$-pVTZ 1.815 0.423 118.834 7.177
$c$-pVQZ 0.279 0.153 0.116 0.565
MgF $c$-pVDZ 0.213 0.122 0.128 1.927
$c$-pVTZ 0.940 2.241 0.180 0.381
$c$-pVQZ 0.312 0.007 0.089 0.116
CaF $c$-pVDZ 0.153 0.090 1.169 0.998
$c$-pVTZ 0.122 0.032 0.264 0.128
$c$-pVQZ 0.086 0.131 0.148 0.099
SrF dyall.v2z 0.037 0.031 0.154 0.341
dyall.v3z 0.164 0.091 1.469 0.085
dyall.v4z 0.180 0.164 1.354 0.006
BaF dyall.v2z 0.024 0.005 0.017 0.340
dyall.v3z 0.011 0.006 0.221 0.027
dyall.v4z 0.195 0.052 0.004 0.118
In summary, we have demonstrated the efficacy of QAE algorithm in computing the PDMs of molecules. Unlike the traditional applications of QAE, which focus on finding the ground-state energy, our approach integrates perturbative techniques and the FFM to enable the calculation of molecular properties. This novel framework shows promise for a broad range of many-body physics applications. Key contributions of our work include: (i) The integration of the finite-field approach into the QAE workflow, (ii) An improved optimization strategy for the Lagrange multipliers within the energy functional, (iii) Perturbative selection and prioritization of important CI coefficients, and (iv) The division of the full QUBO into smaller, independent sub-QUBOs. These sub-QUBOs can be annealed in parallel and subsequently merged for post-processing to recover the final configuration.
The calculated PDM values for molecules like MgF, CaF, SrF, and BaF exhibit excellent agreement with classical benchmark methods, including FCI and RCCSD results, with deviations often below 1%. The results are in agreement with those obtained from SA. The accuracy of the QAE algorithm is rooted in its ability to determine ground-state energies with remarkable precision, as shown in Tables 5 and 6, where absolute errors relative to FCI are consistently on the order of $10^{-5}\%$ or less. The analysis also shows a few anomalies in the PDM calculation. For the BeF molecule in the larger (14o, 7e) active space, QAE performs less accurately with some basis sets, possibly due to how the molecule’s electronic structure interacts with the sub-QUBO partitioning. Aside from these few cases, the method works well across different molecules and basis sets.
Importantly, this work represents the first demonstration of molecular property calculations using quantum annealers. Our findings indicate that QAE, when combined with finite-field methods and perturbative many electron basis selection, can serve as a reliable tool for evaluating electronic properties of molecules. Looking ahead, scaling up QAE to larger active spaces presents challenges due to increased encoding complexity and convergence [37, 38]. However, recent advances in quantum-classical workflows can treat an individual orbital as a qubit for example, varialtional quantum annealing(varQA) [39]. Future advances in quantum annealing hardware, such as new generation D-Wave Advantage2 QPU with Zephyr topology [35] and non-linear quadratic solver [40], may further reduce embedding overheads and improve solution accuracy. These developments are expected to enhance both the QUBO encoding and annealing stages of QAE. We believe that, with continued improvements in hardware and algorithm design—including better preprocessing and error mitigation techniques [34, 41]—quantum annealers may soon play a significant role in atomic and molecular computations, enabling studies beyond the capabilities of classical methods.
Acknowledgements.
Acknowledgment: PPS acknowledges Vikrant Kumar and Aashna Anil Zade for useful discussions on the D-Wave quantum annealer, and Palak Chawla for discussions related to the DIRAC22 package.
Appendix A Appendix A: Ground-State Energy Benchmark Data
Table 5: Ground-state energy comparison (in Hartree) for the (8o, 3e) active space under an applied electric field perturbation of $\epsilon=\pm 0.001$ a.u. QAE and SA results are compared to the benchmark FCI energies, with absolute errors reported.
Molecule Basis FCI SA QAE % Error in QAE
$\epsilon=-0.001$ $\epsilon=+0.001$ $\epsilon=-0.001$ $\epsilon=+0.001$ $\epsilon=-0.001$ $\epsilon=+0.001$ $\epsilon=-0.001$ $\epsilon=+0.001$
BeF $c$-pVDZ -114.218092 -114.198572 -114.218082 -114.198571 -114.218075 -114.198565 1.45E-05 6.19E-06
$c$-pVTZ -114.263250 -114.243655 -114.263246 -114.243646 -114.274282 -114.254704 9.65E-03 9.67E-03
$c$-pVQZ -114.274286 -114.254711 -114.274283 -114.254709 -114.274282 -114.254704 3.60E-06 6.04E-06
MgF $c$-pVDZ -299.548212 -299.471350 -299.548205 -299.471345 -299.548208 -299.471340 1.45E-06 3.23E-06
$c$-pVTZ -299.590767 -299.513813 -299.590766 -299.513812 -299.590766 -299.513812 4.59E-07 3.29E-07
$c$-pVQZ -299.602748 -299.525817 -299.602744 -299.525813 -299.602739 -299.525816 3.04E-06 5.05E-07
CaF $c$-pVDZ -779.318357 -779.171788 -779.318355 -779.171784 -779.318355 -779.171783 2.12E-07 6.28E-07
$c$-pVTZ -779.370339 -779.223768 -779.370338 -779.223768 -779.370336 -779.223768 3.87E-07 5.64E-08
$c$-pVQZ -779.382993 -779.236420 -779.382990 -779.236415 -779.382990 -779.236416 3.31E-07 5.64E-07
SrF dyall.v2z -3277.855874 -3277.560031 -3277.855871 -3277.560027 -3277.855871 -3277.560027 9.57E-08 1.20E-07
dyall.v3z -3277.907353 -3277.611486 -3277.907352 -3277.611483 -3277.907353 -3277.611483 9.60E-09 1.17E-07
dyall.v4z -3277.917885 -3277.622046 -3277.917880 -3277.622044 -3277.917881 -3277.622046 1.22E-07 2.99E-09
BaF dyall.v2z -8235.802103 -8235.346769 -8235.802100 -8235.346766 -8235.802101 -8235.346766 3.28E-08 3.81E-08
dyall.v3z -8235.856833 -8235.401419 -8235.856831 -8235.401417 -8235.856830 -8235.401416 3.51E-08 3.27E-08
dyall.v4z -8235.868050 -8235.412695 -8235.868049 -8235.412693 -8235.868049 -8235.412690 1.79E-08 6.08E-08
Table 6: Ground-state energy comparison (in Hartree) for the (14o, 7e) active space under an applied electric field perturbation of $\epsilon=\pm 0.001$ a.u. QAE and SA results are compared to the benchmark FCI energies, with absolute errors reported.
Molecule Basis FCI SA QAE % error in QAE
$\epsilon=-0.001$ $\epsilon=+0.001$ $\epsilon=-0.001$ $\epsilon=+0.001$ $\epsilon=-0.001$ $\epsilon=+0.001$ $\epsilon=-0.001$ $\epsilon=+0.001$
BeF $c$-pVDZ -114.219629 -114.199386 -114.219477 -114.199272 -114.219306 -114.199230 2.83E-04 1.36E-04
$c$-pVTZ -114.219614 -114.199368 -114.219484 -114.199261 -114.234397 -114.214541 1.29E-02 1.33E-02
$c$-pVQZ -114.274486 -114.254908 114.274458 -114.254887 -114.274449 -114.254870 3.25E-05 3.36E-05
MgF $c$-pVDZ -299.548736 -299.471903 -299.548732 -299.471849 -299.548577 -299.471740 5.31E-05 5.42E-05
$c$-pVTZ -299.590879 -299.513928 -299.590851 -299.513909 -299.590841 -299.513894 1.28E-05 1.13E-05
$c$-pVQZ -299.602824 -299.525899 -299.602798 -299.525876 -299.602797 -299.525870 8.91E-06 9.63E-06
CaF $c$-pVDZ -779.318452 -779.171882 -779.318414 -779.171865 -779.318432 -779.171836 2.63E-06 5.80E-06
$c$-pVTZ -779.370364 -779.223796 -779.370362 -779.223791 -779.370357 -779.223794 9.54E-07 2.37E-07
$c$-pVQZ -779.383009 -779.236437 -779.383007 -779.236433 -779.383006 -779.236431 3.43E-07 7.45E-07
SrF dyall.v2z -3277.888392 -3277.592554 -3277.888370 -3277.592524 -3277.888377 -3277.592535 4.60E-07 5.62E-07
dyall.v3z -3277.939849 -3277.643987 -3277.939839 -3277.643975 -3277.939844 -3277.643978 1.44E-07 2.76E-07
dyall.v4z -3277.950378 -3277.654541 -3277.950373 -3277.654536 -3277.950371 -3277.654541 1.97E-07 8.27E-11
BaF dyall.v2z -8236.102952 -8235.647617 -8236.102945 -8235.647603 -8236.102943 -8235.647608 1.12E-07 1.15E-07
dyall.v3z -8236.157661 -8235.702248 -8236.157655 -8235.702243 -8236.157653 -8235.702236 9.86E-08 1.46E-07
dyall.v4z -8236.168872 -8235.713518 -8236.168870 -8235.713514 -8236.168867 -8235.713513 6.17E-08 6.08E-08
[Uncaptioned image] [Uncaptioned image]
BeF (8o,3e) BeF (14o,7e)
[Uncaptioned image] [Uncaptioned image]
MgF (8o,3e) MgF (14o,7e)
[Uncaptioned image] [Uncaptioned image]
CaF (8o,3e) CaF (14o,7e)
Refer to caption Refer to caption
SrF (8o,3e) SrF (14o,7e)
Refer to caption Refer to caption
BaF (8o,3e) BaF (14o,7e)
Figure 2: Comparison of ground-state energies for molecules under electric field perturbations of $\epsilon=\pm 0.001$ a.u. The charts compare the raw output from the annealer (QAE energy), the energy after classical refinement (QAE-postprocessed energy), and the benchmark FCI energy. The left column shows results for the $(8o,3e)$ active space, and the right column shows results for the $(14o,7e)$ active space.

### v2

@@
 This work provides a pathway for the computation of molecular properties in the quantum annealing paradigm.
-Quantum annealing (QA) is a heuristic approach for solving combinatorial optimization problems that leverages quantum effects such as tunneling and superposition [1].
+Quantum annealing (QA) is a heuristic method for solving combinatorial optimization problems that leverages quantum effects such as tunneling and superposition [1].
 It represents a specialized approach closely related to Adiabatic Quantum Computing, designed to address computationally hard problems by harnessing these quantum effects [2].
-The Quantum Annealer Eigensolver (QAE), a quantum-classical hybrid approach based on the variational principle [3], has been demonstrated to yield highly accurate results for ground-state energy optimization [4].
-Accurate calculation of physical properties is crucial in chemistry and physics as it provides insight to fundamental properties of molecules and materials.
-This method is broadly applicable to a wide range of physical optimization problems, including molecular vibrational spectra [5], complex eigenvalue problems, energy calculations for molecular electronic states, relativistic fine-structure splitting in highly charged atomic ions [4], particle physics simulations [6], and lattice gauge theories [7].
-In this work, we extend the application of QA to compute an important molecular property- the electric permanent dipole moment (PDM).
-This advancement not only underscores the broad applicability of the method but also opens new avenues for analyzing molecular properties beyond ground-state energy evaluation.
+The Quantum Annealer Eigensolver (QAE), a quantum-classical hybrid algorithm based on the variational principle [3], has been demonstrated to yield highly accurate results for ground-state energy optimization [4].
+Accurate calculations of physical properties is important in physics and chemistry as they provide insights to a wide range of applications in molecules and materials.
+The QAE is broadly applicable to a wide range of physical optimization problems, including molecular vibrational spectra [5], complex eigenvalue problems, energy calculations for molecular electronic states, relativistic fine-structure splitting in highly charged atomic ions [4], particle physics simulations [6], and lattice gauge theories [7].
+In this work, we extend the application of QA to compute an important molecular property- the permanent electric dipole moment (PDM).
+This study underscores the broad applicability of the method to molecules, thereby opening a new avenue for computing molecular properties beyond the ground-state energy.
 We compute the PDM of molecules adapting the finite-field method (FFM) to QA.
@@
 The paper concludes with a summary of our key findings and a discussion of potential directions for future research.
-[…]
+The PDM of a molecule is a fundamental property that quantifies the separation of positive and negative charges.
+The FFM offers a numerical technique to evaluate the PDM by measuring the variation of the molecular energy as a response to an external electric field [11, 13].
+When an external electric field $\vec{E}$ is applied along the $z$-direction, the molecular Hamiltonian is perturbed as $\hat{H}=\hat{H}_{0}+\epsilon\hat{O}$ where $\hat{H}_{0}$ is the unperturbed molecular Hamiltonian and $\hat{O}$ is the dipole moment operator in the $z$-direction, and $\epsilon$ is the perturbation strength.
+From first-order perturbation theory, the energy correction due to the field is given by: $E(\epsilon)=E_{0}+\epsilon\left\langle\Psi_{0}\right|\hat{O}\left|\Psi_{0}\right\rangle,$ where $E_{0}$ is the unperturbed energy and $\left|\Psi_{0}\right\rangle$ is the unperturbed wavefunction.
+Thus, the dipole moment in the $z$-direction can be expressed as $\langle\hat{O}\rangle=\left|\frac{\partial E}{\partial\epsilon}\right|_{\epsilon\to 0}$, instead of computing the derivative analytically, we approximate it numerically using the central difference formula:
 $\langle\hat{O}\rangle\approx\lim_{\epsilon\to 0}\frac{E(+\epsilon)-E(-\epsilon)}{2\epsilon}.$ (1)
@@
 Both $E(+\epsilon)$ and $E(-\epsilon)$ are computed using the QAE.
-The Hamiltonian that employed in this work is the DC Hamiltonian, $H_{DC}$, which is:
-$H_{DC}=\sum_{i}\left[c{\alpha}\cdot{p}_{i}+\beta mc^{2}-\sum_{A}\frac{Z_{A}}{|{r}_{i}-{R}_{A}|}\right]+\sum_{i\neq j}\frac{1}{|{r}_{i}-{r}_{j}|}.$ (2)
+The Hamiltonian in this work is the DC molecular Hamiltonian, $\hat{H}_{DC}$, which is:
+$\hat{H}_{DC}=\sum_{i}\left[c{\alpha}\cdot{p}_{i}+\beta mc^{2}-\sum_{A}\frac{Z_{A}}{|{r}_{i}-{R}_{A}|}\right]+\sum_{i\neq j}\frac{1}{|{r}_{i}-{r}_{j}|}.$ (2)
 Here, $c$ is the speed of light, $\alpha$} and $\beta$ are the Dirac matrices, and ${p}_{i}$ refers to the momentum of the $i^{th}$ electron.
@@
 $Z_{A}$ is the atomic number of the $A^{th}$ nucleus.
-Then the second-quantized form of the Hamiltonian is given by
-$H=\sum_{pq}h_{pq}a_{p}^{\dagger}a_{q}+\frac{1}{2}\sum_{pqrs}g_{pqrs}a_{p}^{\dagger}a_{q}^{\dagger}a_{s}a_{r}.$ (3)
-[…]
-The molecular PDMs in this study were first computed using classical computers to provide benchmarks for the values obtained via the QAE algorithm.
+Then, the second-quantized form of the Hamiltonian is given by:
+$\hat{H}=\sum_{pq}h_{pq}a_{p}^{\dagger}a_{q}+\frac{1}{2}\sum_{pqrs}g_{pqrs}a_{p}^{\dagger}a_{q}^{\dagger}a_{s}a_{r}.$ (3)
+where $h_{pq}$ and $g_{pqrs}$ are the one-and two-electron integrals, respectively and $a_{p}^{\dagger}\text{and }a_{q}$ are the fermionic creation and annihilation operators, and the indices run over relativistic orbitals.
+The releativistic equation for molecules can be expressed as: $\hat{H}\left|\Psi\right\rangle=E\left|\Psi\right\rangle$, where $\hat{H}$ is the DC molecular Hamiltonian, $\left|\Psi\right\rangle$ and $E$ are the corresponding wavefunction and energy respectively.
+The wavefunction $\left|\Psi\right\rangle$ is expanded as a linear combination of Slater determinants $\left|\Psi\right\rangle=\sum_{i}c_{i}\left|\Phi_{i}\right\rangle$.
+The determinatal states $\left|\Phi_{i}\right\rangles$ include the DF mean-field and many electron particle-hole states.
+The $c_{i}$ s are the expansion coefficients.
+The molecular PDMs in this study were first calculated using classical computers to provide benchmarks for the values obtained via the QAE algorithm.
 Specifically, DF and relativistic coupled-cluster singles and doubles (RCCSD) calculations were performed using the DIRAC22 software package [14].
 The DF computations were carried out at the self–consistent field (SCF) level, with a perturbative electric field ($\epsilon=\pm 10^{-3}$) applied along the $z$-axis to evaluate the field-dependent energies.
-To enhance computational efficiency, molecular symmetry was exploited through use of C2v double group symmetry in the molecular setup.
-In addition to DF, RCCSD calculations were performed using the same code, employing the one- and two-electron integrals generated internally by DIRAC22.
+To enhance computational efficiency, molecular symmetry was exploited through use of C2v double group symmetry [15].
 To enable a direct comparison with the QAE algorithm, the FCI Hamiltonian was constructed classically by applying the Slater–Condon rules [16] and subsequently diagonalized exactly to obtain the FCI/ED energy.
@@
 The ground-state energies corresponding to both positive and negative perturbation strengths ($\epsilon=\pm 10^{-3}$) were computed to evaluate the field-dependent energy shifts.
-QA was employed to determine the electronic energies within selected active spaces—specifically, $(8o,3e)$ and $(14o,7e)$ configurations.
+QA was employed to determine the electronic energies in active spaces $(No,Me)$ = $(8o,3e)$ and $(14o,7e)$ where, $M$ denotes the number of occupied orbitals out of $N$ total orbitals.
 These electronic energies were then combined with the corresponding core energy and nuclear repulsion energy terms to obtain the total molecular energy.
-Using this total energy, the electronic component of the dipole moment was evaluated through the finite-field approach explicitly linked Eq.(1).
-To obtain the PDM, the nuclear contribution was added based on experimental available bond lengths.
+Using this total energy, the electronic component of the dipole moment was evaluated through the finite-field approach as shown by Eq.(1).
+To obtain the PDM, the nuclear contribution was added based on experimentally available bond lengths.
 The bond lengths used in the calculations were 1.361 Å for BeF, 1.75 Å for MgF, 1.967 Å for CaF, 2.075 Å for SrF, and 2.16 Å for BaF [17, 18, 19].
-For all calculations, uncontracted Gaussian-type orbitals (GTOs) [19] were used, and the kinetic balance condition was strictly maintained to ensure the accuracy of relativistic treatments.
-The basis sets for the heavier atoms Sr and Ba were chosen from the dyall.c2v, dyall.c3v [20], and dyall.c4v [21, 19].
-For lighter elements such as Be, Mg, Ca, and F, the exponential parameters from the correlation-consistent polarized valence (cc-pV) basis sets were used, obtained from the EMSL Basis Set Exchange Library [20].
-Table 1: Atomic basis sets used in the calculations.
-This table lists the uncontracted GTOs employed for each atom.
-For Be, Mg, Ca, and F, correlation-consistent polarized valence (cc-pV) basis sets of increasing size were used.
-For the heavier Sr and Ba atoms, Dyall’s relativistic basis sets (dyall.c2v, dyall.c3v, dyall.c4v) were utilized.
-   Atom Basis
-Be cc-pVDZ: 9s, 4p, 1d
-cc-pVTZ: 11s, 5p, 2d, 1f
-cc-pVQZ: 12s, 6p, 3d, 2f, 1g
-Mg cc-pVDZ: 12s, 8p, 1d
-cc-pVTZ: 15s, 10p, 2d, 1f
-cc-pVQZ: 16s, 12p, 3d, 2f, 1g
-Ca cc-pVDZ: 14s, 11p, 5d
-cc-pVTZ: 20s, 14p, 6d, 1f
-cc-pVQZ: 22s, 16p, 7d, 2f, 1g
-Sr dyall.c2v: 20s, 14p, 9d
-dyall.c3v: 28s, 20p, 13d, 2f
-dyall.c4v: 33s, 25p, 15d, 4f, 2g
-Ba dyall.c2v: 25s, 19p, 13d
-dyall.c3v: 31s, 25p, 15d, 2f
-dyall.c4v: 36s, 30p, 18d
-F cc-pVDZ: 9s, 4p, 1d
-cc-pVTZ: 10s, 5p, 2d, 1f
-cc-pVQZ: 12s, 6p, 3d, 2f, 1g
-QA involves preparing the ground-state of an initial Hamiltonian, $H_{I}$, which is typically chosen to be the transverse field Hamiltonian, and then adiabatically transforming it into another Hamiltonian, $H_{F}$, generally chosen to be the Ising Hamiltonian whose ground-state encodes the solution to the problem.
+For all the calculations, uncontracted Gaussian-type orbitals (GTOs) [20] were used, and the kinetic balance condition was implemented to ensure the proper relationship between the large and small components of the orbitals.
+Each orbital is represented by n-Slater type functions, each with a different value of exponent (zeta) [15].
+The basis sets for the heavier atoms Sr and Ba were chosen from dyall.c2v, dyall.c3v [21], and dyall.c4v [22, 20] For lighter elements such as Be, Mg, Ca, and F, using exponents from the correlation-consistent polarized valence (cc-pV) basis, which were obtained from the EMSL Basis Set Exchange Library [21].
+QA involves preparing the ground-state of an initial Hamiltonian, $\hat{H}_{I}$, which is typically chosen to be the transverse field Hamiltonian, and then adiabatically transforming it into another Hamiltonian, $\hat{H}_{F}$, generally chosen to be the Ising Hamiltonian whose ground-state encodes the solution to the problem.
 The transformation follows:
-$H(t)=\left(1-\frac{t}{T}\right)H_{I}+\frac{t}{T}H_{F};\quad t\in[0,T],$ (4)
-[…]
-$\xi(\vec{c},\lambda)=\sum_{i,j=0}^{N-1}c_{i}c_{j}(H_{ij}-\lambda\delta_{ij}).$ (5)
-[…]
+$\hat{H}(t)=f(t)\hat{H}_{I}+g(t)\hat{H}_{F};\quad t\in[0,T],$ (4)
+where the system evolves smoothly from the ground state of $\hat{H}_{I}$ to that of $\hat{H}_{F}$ during the annealing time $T$.
+Initially, $f(0)\gg g(0)$, giving a simple delocalized ground state with spins $s_{i}=\pm 1$.
+As time progresses, $f$ decreases and $g$ increases, and by $t=T$, $f(T)\ll g(T)$, completing the annealing process.
+While the adiabatic model described above is the theoretical foundation of QA, in practice, quantum annealing operates far from the adiabatic limit.
+In this work, we apply the QAE algorithm to the electronic structure problem.
+The QAE algorithm solves the eigenvalue equation $\hat{H}\left|\Psi\right\rangle=E\left|\Psi\right\rangle$ by transforming it into an energy minimization problem [23, 24].
+The energy functional considered is the expectation value of the Hamiltonian $\hat{H}$ with respect to an unknown state $\left|\Psi\right\rangle$.
+To avoid trivial solutions, the normalization constraint $\left\langle\Psi\middle|\Psi\right\rangle-1=0$ is enforced by including a Lagrange multiplier $\lambda$ into the energy functional.
+In the modified formulation, after dropping irrelevant constants the energy functional will be $\xi=\left\langle\Psi\right|\hat{H}\left|\Psi\right\rangle-\lambda\left\langle\Psi\middle|\Psi\right\rangle$.
+Assuming the wavefunction is expressed as $\left|\Psi\right\rangle=\sum_{i=1}^{N}c_{i}\left|\Phi_{i}\right\rangle$, where $\{\left|\Phi_{i}\right\rangle\}$ stands for a collection of $N$ predefined basis functions (determinatal states), the QAE algorithm seeks to find the coefficients $\{c_{i}\}$, with each $c_{i}$ lying in the interval [1,1][-1,1], for the energy functional:
+$\xi(\vec{c},\lambda)=\sum_{i,j=1}^{N}c_{i}c_{j}(H_{ij}-\lambda\delta_{ij}).$ (5)
+In this context, $H_{ij}\equiv\left\langle\Phi_{i}\right|\hat{H}\left|\Phi_{j}\right\rangle$ and $\delta_{ij}$ denotes the Kronecker delta function.
+To translate the problem for quantum annealing, we implemented fixed-point encoding [5] strategy wherein each coefficient $c_{i}$ is expressed using $K$ qubits $q_{\alpha}^{i}\in\{0,1\}$:
 $c_{i}=\sum_{\alpha=1}^{K-1}2^{\alpha-K}q_{\alpha}^{i}-q_{K}^{i}.$ (6)
 This converts the problem from a continuous optimization over $\{c_{i}\}$ to a discrete optimization over the binary variables $\{q^{i}_{\alpha}\}$.
-Here, the first K1K-1 qubits encode the fractional part, the last qubit $q_{K}^{i}$ encodes the sign and So the each $c_{i}$ is repesented by $K$ number of binary variable $q^{i}_{\alpha}$ [24, 25, 5].
+Here, the first K1K-1 qubits encode the fractional part, the last qubit $q_{K}^{i}$ encodes the sign and thus the each $c_{i}$ is repesented by $K$ number of binary variables $q^{i}_{\alpha}$ [25, 26, 5].
+The energy functional becomes:
 $\begin{split}\xi(\vec{q},\lambda)=\sum_{i,j=1}^{N}\sum_{\alpha,\beta=1}^{K-1}&\left(2^{\alpha+\beta-2K}q_{\alpha}^{i}q_{\beta}^{j}-2^{\alpha+1-K}q_{\alpha}^{i}+q_{K}^{i}q_{K}^{j}\right)\\ &\times\left(H_{ij}-\lambda\delta_{ij}\right).\end{split}$ (7)
 The proposed mapping approach necessitates an expansion of the Quadratic Unconstrained Binary Optimization (QUBO) formulation [27] from dimensions $(N\times N)$ to $(NK\times NK)$.
-This dimensional increase is required to overcome the restricted connectivity limitations inherent in D-Wave’s quantum annealing architecture.
-So, the optimization of $\xi(\vec{q},\lambda)$ is implemented through a pipelined framework consisting of sequential sub-QUBO optimizations, where each sub-problem operates on an independent subset of coefficients.
-In our methodology, we first sort the coefficients in descending order basis on the basis of their classical values obtained from perturbation theory.
-These ordered coefficients are then partitioned into systematic groupings that form the basis for sub-QUBOs construction.
-These sub-QUBOs are then independently submitted to the annealer to obtain corresponding solution [4].
-The embedding process maps the each sub-QUBO onto the D-Wave hardware’s native topology [27, 28] by forming chains of physical qubits to represent single logical variables.
-We used the default chain coupling strength to maintain consistent correlations during annealing while preventing chain dynamics from dominating the system’s energy landscape [29].
-During the annealing process, we first map the subproblem onto the Pegasus graph structure of the D-Wave Advantage quantum processing unit.
-For each execution, we acquire a fixed ensemble of 1000 samples(anneals or number of reads), where each sample represents a specific configuration of the coefficients in the qubit state space.
-After obtaining the binary solutions from each independent sub-QUBO problem from annealer, we aggregate those solutions and proceed to a post-processing stage.
-This stage involves applying a steepest descent optimization at every value of $\lambda$ within the defined search interval $\lambda_{-},\lambda_{+}]$.
-At every value of $\lambda$ the annealer returns binary values of $\vec{q}$}, totaling $NK$ bits.
-In our case, we have set $K=10$ to encode each parameter with sufficient resolution.
-The entire implementation utilizes the tools available in D-Wave’s Ocean SDK [30], and specifically employs the D-Wave Advantage QPU for execution.
-This method begins with an $NK$-variable QUBO problem, due to the computational complexity of directly solving such large-scale problems, we employ a perturbative approximation to reduce the problem size while preserving the essential features of the solution space.
-In this work, the strategic partitioning of a large-scale QUBO problem into $M$ smaller, more tractable sub-QUBOs is employed.
+In view of restricted connectivity limitations inherent to D-Wave’s quantum annealing architecture, the optimization of $\xi(\vec{q},\lambda)$ is implemented through a pipelined framework consisting of sequential $M$ smaller, more tractable sub-QUBO optimizations, where each sub-problem of energy functional $\xi_{m}(\vec{q}^{(m)},\lambda)$ operates on an independent subset of variables.
 Each resulting sub-problem is then optimized independently to find a local solution, denoted as $\vec{q}^{(m)}$ for $m\in\{1,\dots,M\}$.
@@
 Although the decomposition methodology permits parallel execution on multiple quantum annealers, in this work the sub-QUBOs are optimized sequentially due to limited processor access, and thus the overall computational demand remains unchanged.
-Figure 1: Workflow of the standard QAE algorithm, where an energy functional with Lagrange multiplier $\lambda$ is mapped to a QUBO, optimized on the quantum annealer, and refined via classical post-processing across a range of $\lambda$ values to obtain the minimum ground-state energy.
-The global solution is constructed by combining partial solutions into a candidate vector $\vec{q}$.
+In our methodology, we first sort the coefficients in descending order on the basis of their values obtained classically from perturbation theory.
+These ordered coefficients are then partitioned into systematic groupings that form the basis for sub-QUBOs construction to reduce the problem size while preserving the essential features of the solution space.
+These sub-QUBOs are then independently submitted to the annealer to obtain the corresponding solution [4].
+The embedding process maps the each sub-QUBO onto the D-Wave hardware’s native topology [28, 29] by forming chains of physical qubits to represent single logical variables.
+We used the default chain coupling strength during annealing [30].
+During the annealing process, we first map the subproblem onto the Pegasus graph structure of the D-Wave Advantage quantum processing unit (QPU).
+For each execution, we acquire a fixed ensemble of 1000 samples(anneals or number of reads), where each sample represents a specific configuration of the coefficients in the qubit state space.
+After obtaining the binary solutions from each independent sub-QUBO problem from the quantum annealer, we aggregate those solutions and proceed to a post-processing stage.
+This stage involves applying a steepest descent optimization at every value of $\lambda$ within the defined search interval $\lambda_{\text{min}},\lambda_{\text{max}}]$.
+For each value of $\lambda$ the annealer returns binary values of $\vec{q}^{(m)}_{\text{opt}}$} for each sub-QUBO.
+In our case, we have set $K=10$ to encode each parameter with sufficient resolution.
+The entire implementation utilizes the tools available in D-Wave’s Ocean SDK [31], and specifically employs the D-Wave Advantage quantum processing units(QPU) for execution.
+Then, the global solution $(\vec{q}_{\text{raw}})$ of size $NK$ variables is constructed by combining partial solutions $({\vec{q}^{(m)}_{\text{opt}}})$.
 We employ steepest descent minimization within the energy landscape $\xi(\vec{q},\lambda)$, where at each iteration the algorithm selects the direction of maximal energy reduction.
-This refinement stage compensates for decomposition approximations by explicitly incorporating all intervariable couplings from the original Hamiltonian.
-The converged solution $\vec{q}_{\text{final}}$ attains accuracy comparable to direct methods.
-Algorithm 1 QA for Hamiltonian Minimization
+This refinement stage compensates for decomposition approximations by explicitly incorporating all inter-connected couplings from the original Hamiltonian.
+The converged solution $\vec{q}_{\text{final}}$ attains accuracy comparable to classical methods.
+Algorithm 1 QA for Energy Minimization
 1:Full Hamiltonian matrix $H$ of size $N\times N$.
@@
 13:Phase 3: Post-processing and Energy Evaluation
-14:  Apply a steepest descent method for refinement: $\vec{q}_{\text{post-processed}}\leftarrow\text{SteepestDescent}(\vec{q}_{\text{raw}})$.
-15:  Reverse the encoding to obtain classical coefficients: $\vec{c}_{\text{post-processed}}\leftarrow\text{ReverseEncoding}(\vec{q}_{\text{post-processed}})$.
-16:  Compute the total energy $E_{\lambda}$ using the obtained coefficients: $E_{\lambda}\leftarrow\text{ComputeEnergy}(\vec{c}_{\text{post-processed}})$.
+14:  Apply a steepest descent method for refinement: $\vec{q}_{\text{final}}\leftarrow\text{SteepestDescent}(\vec{q}_{\text{raw}})$.
+15:  Reverse the encoding to obtain classical coefficients: $\vec{c}_{\text{final}}\leftarrow\text{ReverseEncoding}(\vec{q}_{\text{final}})$.
+16:  Compute the total energy $E_{\lambda}$ using the obtained coefficients: $E_{\lambda}\leftarrow\text{ComputeEnergy}(\vec{c}_{\text{final}})$.
 17:end for
@@
 20:Return $\vec{c}_{\text{final}}$
-In Table 2 and 3, we present our computed PDM values using the QAE and benchmark them against conventional methods such as RCCSD, FCI, and SA results [31], all within two different active spaces but employing different basis sets as described in Table 1.
+Construct sub-QUBOs sub-QUBO energy functional $\xi_{m}(\vec{q}^{(m)},\lambda)$ Optimization Embedding & annealing on QPU + postprocessing Wavefunction construction Build $|\Psi\rangle$ from qubit configuration Energy evaluation $\langle\Psi|\hat{H}|\Psi\rangle$ $\lambda\in[\lambda_{\min},\lambda_{\max}]$
+Figure 1: Workflow of the QAE algorithm, where an energy functional with Lagrange multiplier $\lambda$ is mapped to a sub-QUBO, optimized on the quantum annealer, and refined via post-processing across $\lambda$ values to obtain the minimum ground-state energy.
+In Tables 1 and 2, we present our computed PDM values using the QAE and benchmark them against conventional methods such as RCCSD, FCI, and SA results [32], all within two different active spaces but employing different basis sets.
 The QAE computations were performed on both classical (SA) platforms [32] and quantum annealing hardware (D-Wave Advantage).
-These simulations explicitly take into account qubit connectivity constraints.
+These simulations explicitly took into account qubit connectivity constraints.
 The results indicate that QAE can accurately predict both total energies (for both perturbation $\pm\epsilon$) and PDM values, with deviations within acceptable limits.
 Specifically, the hardware-executed QAE results were found to be consistent with their SA and FCI counterparts as given in Table 3.
-For the QAE calculations, we selected energy ranges $\lambda_{-},\lambda_{+}]$ slightly below and above the DF energy, corresponding to the active spaces $(8o,3e)$ and $(14o,7e)$ used in this study.
+For the QAE calculations, we selected energy ranges $\lambda_{\text{min}},\lambda_{\text{max}}]$ slightly below and above the DF energy, corresponding to the active spaces $(8o,3e)$ and $(14o,7e)$ used in this study.
 After generating the full configuration interaction determinant basis, we employed a perturbative selection procedure to retain only the most relevant configurations.
@@
 The Hamiltonian matrix was then constructed within the reduced determinant space and used for subsequent QAE calculations.
-For each specified energy range $\lambda_{-},\lambda_{+}]$, we performed three independent QAE runs which called repeats and selected the lowest energy [32, 5].
+For each specified energy range $\lambda_{\text{min}},\lambda_{\text{max}}]$, we performed three independent QAE runs which we term as repeats and selected the lowest energy [33, 5].
 Additionally, sub-QUBOs were restricted to a maximum of 60 binary variables for practical implementation on the hardware in this calculation.
@@
 Furthermore, we conducted simulations assuming all-to-all connectivity and using the same sub-QUBO decomposition as in the quantum hardware setup.
-This residual discrepancy is primarily attributed to limitations in our QUBO encoding scheme—specifically, the choice of $K=10$ in the fixed-point representation of the CI coefficients, yielding a precision of $\frac{1}{2^{10}}$ in energy.
+This residual discrepancy is primarily attributed to limitations in our QUBO encoding scheme—specifically, the choice of $K=10$ in the fixed-point representation of the expansion coefficients, yielding a precision of $\frac{1}{2^{10}}$ in energy.
 Increasing $K$ would improve numerical precision but also significantly enlarge the QUBO, rendering it less feasible for hardware execution at present.
 The most prominent deviation was observed between simulations assuming all-to-all connectivity [36] and those constrained by the native connectivity of the D-Wave hardware, indicating that chain formation and embedding limitations are key sources of inaccuracy [30].
-Finally, we note that our method does not suffer from barren plateaus, which are common in approaches like VQE [36] due to the structure of the ansatz.
-In contrast, QAE directly optimizes the CI coefficients, and the energy remains sensitive to changes in these values.
-Since our approach is based on complete active space CI, it avoids the flat regions in the energy landscape that make optimization difficult.
-Moreover, we were able to successfully calculate properties such as the PDM using this method, demonstrating its robustness compared to VQE.
-Table 2: Comparison of the PDM values, in debye, for the $(8o,3e)$ and $(14o,7e)$ active space.
+Finally, we note that our method does not seems to suffer from barren plateau and local minima, which are common in approaches like VQE [37, 38].
+In QAE, the energy functional is optimized to obtain the qubit binary variables, from which the expansion coefficients are subsequently determined, as shown in Eq.
+(6).
+Since this approach does not involve gate-based operations, it is free from gate errors that typically affect the performance of variational algorithms like VQE.
+Moreover, for the SrF molecule in the $(14o,7e)$ active space, our method achieved an accuracy of about $98.6\%$, while comparable accuracy in VQE would require extensive circuit optimization and error mitigation techniques as exemplified by a 6-qubit VQE computation on IonQ hardware, where an accuracy of $93.1\%$ was achieved [39].
+The accurate computation of properties such as the PDM further demonstrates the robustness and reliability of QAE compared to VQE.
+Table 1: Comparison of the PDM values, in Debye, for the $(8o,3e)$ active space.
 Results from the QAE are benchmarked against classical RCCSD, FCI, and SA methods for the molecules and basis sets studied.
@@
 Table 3: Percentage error for the PDM values calculated via the QAE and SA.
-The error is calculated relative to the benchmark FCI results presented in Tables II and III for the $(8o,3e)$ and $(14o,7e)$ active spaces, respectively
+The error is calculated relative to the benchmark FCI results presented in Tables II and III for the $(8o,3e)$ and $(14o,7e)$ active spaces, respectively.
 Molecule Basis (8o,3e) (14o,7e)
-$Err_{QAE}$ $Err_{SA}$ $Err_{QAE}$ $Err_{SA}$
+$\mathrm{Err_{QAE}}$ $\mathrm{Err_{SA}}$ $\mathrm{Err_{QAE}}$ $\mathrm{Err_{SA}}$
 BeF $c$-pVDZ 0.894 0.693 50.837 11.738
@@
 Unlike the traditional applications of QAE, which focus on finding the ground-state energy, our approach integrates perturbative techniques and the FFM to enable the calculation of molecular properties.
-This novel framework shows promise for a broad range of many-body physics applications.
-Key contributions of our work include: (i) The integration of the finite-field approach into the QAE workflow, (ii) An improved optimization strategy for the Lagrange multipliers within the energy functional, (iii) Perturbative selection and prioritization of important CI coefficients, and (iv) The division of the full QUBO into smaller, independent sub-QUBOs.
+This novel framework holds promise for a broad range of many-body physics applications.
+Key contributions of our work include the integration of the finite-field approach into the QAE workflow and the division of the full QUBO into smaller, independent sub-QUBOs by selecting important expansion coefficients.
 These sub-QUBOs can be annealed in parallel and subsequently merged for post-processing to recover the final configuration.
-The calculated PDM values for molecules like MgF, CaF, SrF, and BaF exhibit excellent agreement with classical benchmark methods, including FCI and RCCSD results, with deviations often below 1%.
+The calculated PDM values for MgF, CaF, SrF, and BaF exhibit excellent agreement with classical benchmark methods, including FCI and RCCSD results, with deviations of average below 1%.
 The results are in agreement with those obtained from SA.
-The accuracy of the QAE algorithm is rooted in its ability to determine ground-state energies with remarkable precision, as shown in Tables 5 and 6, where absolute errors relative to FCI are consistently on the order of $10^{-5}\%$ or less.
+An important feature of relativistic coupled-cluster method is that it includes physical effects to all orders of perturbation in the residual Coulomb interaction for all n particle-n hole excitations.
+Tables I and II of supplementary material indicate that absolute errors relative to FCI are consistently on the order of $10^{-5}\%$ or less.
 The analysis also shows a few anomalies in the PDM calculation.
@@
 Aside from these few cases, the method works well across different molecules and basis sets.
-Importantly, this work represents the first demonstration of molecular property calculations using quantum annealers.
-Our findings indicate that QAE, when combined with finite-field methods and perturbative many electron basis selection, can serve as a reliable tool for evaluating electronic properties of molecules.
-Looking ahead, scaling up QAE to larger active spaces presents challenges due to increased encoding complexity and convergence [37, 38].
-However, recent advances in quantum-classical workflows can treat an individual orbital as a qubit for example, varialtional quantum annealing(varQA) [39].
-Future advances in quantum annealing hardware, such as new generation D-Wave Advantage2 QPU with Zephyr topology [35] and non-linear quadratic solver [40], may further reduce embedding overheads and improve solution accuracy.
+Importantly, this work represents the first computation of a molecular property other than energy using a quantum annealer.
+Our findings indicate that QAE, when combined with finite-field methods and perturbative many-electron basis selection, could serve as a reliable tool for evaluating electronic properties of molecules.
+Scaling up QAE to larger active spaces presents challenges due to increased encoding complexity and subtle convergence issues [40, 41].
+However, recent advances in quantum-classical workflows can treat an individual orbital as a qubit as in, variational quantum annealing (varQA) algorithm [42].
+Future advances in quantum annealing hardware, such as new D-Wave Advantage2 QPU with Zephyr topology [36] and non-linear quadratic solver [43], may further reduce embedding overheads and improve solution accuracy.
 These developments are expected to enhance both the QUBO encoding and annealing stages of QAE.
-We believe that, with continued improvements in hardware and algorithm design—including better preprocessing and error mitigation techniques [34, 41]—quantum annealers may soon play a significant role in atomic and molecular computations, enabling studies beyond the capabilities of classical methods.
+Quantum annealers may soon play a significant role in atomic and molecular computations, enabling studies beyond the capabilities of classical methods with continued improvements in hardware and algorithm design—including better preprocessing and error mitigation techniques [35, 44].
 Acknowledgements.
 Acknowledgment: PPS acknowledges Vikrant Kumar and Aashna Anil Zade for useful discussions on the D-Wave quantum annealer, and Palak Chawla for discussions related to the DIRAC22 package.
-Appendix A Appendix A: Ground-State Energy Benchmark Data
+Appendix A SUPPLEMENTARY MATERIAL : Ground-State Energy Benchmark Data
 Table 4: Ground-state energy comparison (in Hartree) for the (8o, 3e) active space under an applied electric field perturbation of $\epsilon=\pm 0.001$ a.u.
@@
 BeF (8o,3e) BeF (14o,7e)
-[Uncaptioned image] [Uncaptioned image]
+Refer to caption Refer to caption
 MgF (8o,3e) MgF (14o,7e)
-[Uncaptioned image] [Uncaptioned image]
+Refer to caption Refer to caption
 CaF (8o,3e) CaF (14o,7e)